Could someone please help me solve the system by graphing then classify the system as consistent or inconsistent and dependent or independent?
8x-5y=20
5y-8x=-20

Answer by ewatrrr(24785)

 
Hi
solve the system by graphing
8x-5y=20 (0,-4) and (20/8,0)
5y-8x=-20 ( 0,-4) and (20/8,0)
Graphing points and connecting with a line:
These are the same lines, infinite solutions. System is Dependent.
A consistent system has at least one solution.
An inconsistent system has no solutions.
An independent system has exactly one solution.
A dependent system has infinitely many solutions.